0 0
Add a comment Improve this question Transcribed image text
Answer #1

The true far point should be infinity, so we are taking objects from that distance and focusing them at 4.50 m

Apply 1/f = 1/p + 1/q

1/f = 1/-4.5

f = -4.5 m

Now, for the new near point...

1/-4.5 = 1/.24 + 1/q

q = -22.8

Thus the new near point is 22.8 cm